OpenStudy (anonymous):

Write the equation of the line that passes through (−1, 5) and has a slope of 3 in point-slope form. (2 points) x + 1 = 3(y − 5) x − 5 = 3(y + 1) y − 5 = 3(x + 1) y + 1 = 3(x − 5)

OpenStudy (ayee_ciera):

y-y1=m(x-x1)

OpenStudy (ayee_ciera):

plug in your numbers

OpenStudy (teddyiswatshecallsme):

as @aye_ciera said, we're going to use the point slope formula to write this equation. m = slope x1 = first point y1 = second point so our m is 3 and our ordered pair is (-1, 5) so all we have to do is rewrite the equation with the points in it(and #'s) y - (5) = 3(x + 1) all we did is plug in the numbers for our equation. So you're answer would be C. :3